# DeepSafeMPC: Deep Learning-Based Model Predictive Control for Safe Multi-Agent Reinforcement Learning

### 11 Mar 2024 v1

@@ IV-A Multi-Agent Proximal Policy Optimization
$+\sigma\left[\frac{1}{Bn}\sum_{i=1}^{B}\sum_{t=1}^{n}S[\pi_{\theta}(o_{i}^{t})]\right].$
where $r_{\theta}^{t}=\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})},\ A_{i}^{t}$ is computed using the GAE [28] method, $S$ is the policy entropy, $\epsilon$ is the clip parameter and $\sigma$ is the entropy coefficient hyperparameter.
Certainly, here’s a polished version of your statement regarding centralized critic networks in the context of Multi-Agent Proximal Policy Optimization (MAPPO):
In contrast, centralized critic networks aim to update their value function estimations by minimizing a carefully designed loss function. This function quantifies the discrepancy between estimated returns and actual observed returns, leveraging global information to achieve a comprehensive understanding of the environment. By assessing the collective information, the centralized critic provides a crucial feedback mechanism. This mechanism guides individual agents in optimizing their policies, ensuring that each agent’s decisions are informed by not just their own experiences, but also by the collective experiences of all agents within the system. Such an approach is instrumental in addressing the inherent non-stationarity of multi-agent systems.
@@
In these formulations, $B$ represents the batch size and $n$ the number of agents.
@@ IV-B Dynamics Predictor
During the training, we use an offline scheme to learn the set of parameters $\eta$. We gather extensive operational data states $s^{t}$, actions $\mathbf{a}^{t}$ and next states ${s}^{t+1}$ from the environment, which is subsequently buffered. Then we use ${s}^{t+1}$ as the ground truth and minimize the error between the ground truth and generated next states $\hat{s}^{t+1}$. For the optimization objective, we leverage the Mean Squared Error (MSE) loss as follows:
@@
By leveraging historical data, we can construct a dataset that accurately captures the transitions between states under various action conditions. This dataset enables a neural network to learn the temporal dynamics governing the evolution of states within the environment. To ascertain the robustness of the MPC’s predictor, we provide a straightforward proof and demonstrate the prediction error through experimental results.
In the robustness analysis, our primary objective is to prove that the training error of the proposed method decreases incrementally over time. The training error is denoted by $e^{t}=\hat{s}^{t}-s^{t}$, where $\hat{s}^{t}$ and $s^{t}$ are defined as follows:
$h_{1}(s^{t},\mathbf{a}^{t})=W^{*}\eta(s^{t},\mathbf{a}^{t}),$ (8)
$h_{2}(s^{t},\mathbf{a}^{t})=W\eta(s^{t},\mathbf{a}^{t}).$ (9)
Here, $h_{1}(\cdot)$ and $h_{2}(\cdot)$ represent nonlinear functions, $W^{*}$ denotes the optimal neural network weights, while $W$ represents the actual neural network weights during training, and $\sigma(\cdot)$ represents the activation function. Our theoretical proof introduces the assumption:
@@
The training samples comprise bounded sequences, the neural network’s weights $W$ are bounded, and there exists a set of optimal weights $W^{*}$.
@@
Assuming the validity of Assumption 1, we have:
$\|W^{*}-W\|<\alpha,$ (10)
$\|h_{1}(s^{t},a^{t})-h_{2}(s^{t},a^{t})\|=\|e(t)\|\leq\varepsilon_{f},$ (11)
@@
Given that $W^{*}$ is constant and $W$ is bounded, a positive real number $\alpha$ exists such that $\|W^{*}-W\|<\alpha$. Additionally, to mitigate the gradient exploding problem, the network’s output is typically constrained within a specific range, as represented by Equation (11). ∎
@@
With the deep learning-based predictive model in place, the training error $e(t)$ is uniformly ultimately bounded (UUB) according to Lemma IV.1.
@@
We employ the Lyapunov function to analyze prediction errors:
@@
$=e^{T}(t)(h_{1}(s^{t},\mathbf{a}^{t})-h_{2}(s^{t},\mathbf{a}^{t})-e(t))$
$\leq\|e(t)\|^{2}+\|e(t)\|\|W^{*}\eta(s^{t},\mathbf{a}^{t})-W\eta(s^{t},\mathbf{a}^{t})\|$
$=\|e(t)\|^{2}+\|e(t)\|\|\eta(s^{t},\mathbf{a}^{t})\|\|W^{*}-W\|$
$\leq\|e(t)\|^{2}+\|e(t)\|\|\eta(s^{t},\mathbf{a}^{t})\|\alpha.$
@@
  1. 1.
    Where the error norm $\|e(t)\|$ meets or exceeds a threshold $\alpha$, the derivative of the Lyapunov function $\dot{V}(e(t))$ does not surpass zero. Concurrently, the error norm is bounded above by $e_{f}$, thereby ensuring that the training error remains confined within a predetermined range.
  2. 2.
    When the error norm $\|e(t)\|$ falls short of $\alpha$, the derivative of the Lyapunov function becomes positive, indicating an escalation in the training error up to the point where $\|e(t)\|$ reaches the threshold $\alpha$. If the error norm equals or exceeds $\alpha$, then $\dot{V}(e(t))$ will become non-positive, leading to an error norm that is ultimately bounded by $e_{f}$.
In light of these findings, it can be concluded that the training error will asymptotically be less than or equal to $e_{f}$, which proves the robustness of the predictor model. ∎
@@ IV-C Integration with MPC
By leveraging the deep learning-based predictor, we integrate it into MPC to deal with the implicit dynamics. The goal of MPC in this work is to find a suitable control signal $\mathbf{a}^{t}$ through online optimization. In this work, given prediction and control horizons $T$, the optimization problem of MPC can be formulated as follows:
@@
$+\sum_{\tau=t+1}^{t+T}\lambda_{\tau}\left(f\left(s^{\tau},\mathbf{a}^{\tau}\right)-\hat{y}^{\tau}\right)$
@@
In SQP, each iteration involves solving a QP subproblem to obtain a search direction for the optimization variables. The QP subproblem is formulated based on a quadratic approximation of the objective function and a linear approximation of the constraints around the current iterate $k$. And the QP subproblem for a particular iteration $(s^{k},\mathbf{a}^{k})$ can be stated as follows:

### 12 Mar 2024 v2

@@ IV-A Multi-Agent Proximal Policy Optimization
$+\sigma\left[\frac{1}{Bn}\sum_{i=1}^{B}\sum_{t=1}^{n}Z[\pi_{\theta}(o_{i}^{t})]\right].$
where $r_{\theta}^{t}=\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})},\ A_{i}^{t}$ is computed using the GAE [28] method, $\epsilon$ is the clip parameter, $\sigma$ is the entropy coefficient hyperparameter and $Z$ is the policy entropy, which encourages the exploration of new actions.
In contrast, centralized critic networks aim to update their value function estimations by minimizing a carefully designed loss function. This function quantifies the discrepancy between estimated returns and actual observed returns, leveraging global information to achieve a comprehensive understanding of the environment. By assessing the collective information, the centralized critic provides a crucial feedback mechanism. This mechanism guides individual agents in optimizing their policies, ensuring that each agent’s decisions are informed by not just their own experiences, but also by the collective experiences of all agents within the system.
@@
In these formulations, $B$ represents the batch size and $n$ the number of agents. This detailed exploration of MAPPO highlights its sophisticated approach to multi-agent reinforcement learning, showcasing the critical roles of policy entropy and centralized critique in developing effective, adaptive cooperative strategies.
@@ IV-B Dynamics Predictor
During the training, we use an offline scheme to learn the set of parameters $\eta$. We gather extensive operational data states $s^{t}$, actions $\mathbf{a}^{t}$ and next states ${s}^{t+1}$ from the environment, which is subsequently buffered. Then we leverage ${s}^{t+1}$ as the ground truth and minimize the error between the ground truth and generated next states $\hat{s}^{t+1}$. For the optimization objective, we utilize the Mean Squared Error (MSE) loss as follows:
@@
By leveraging historical data, we construct a dataset that accurately captures the transitions between states under a range of action conditions. This dataset enables a neural network to learn the temporal dynamics governing the evolution of states within the environment. To ascertain the robustness of the MPC’s predictor, we provide a straightforward proof and demonstrate the prediction error through experimental results.
In the robustness analysis, our primary objective is to prove that the training error of the proposed method diminishes progressively over time. The training error is denoted by $e^{t}=\hat{s}^{t}-s^{t}$, and we redefine the equation (2) as follows:
$f_{1}(s^{t},\mathbf{a}^{t})=\sigma(\eta^{*}(s^{t},\mathbf{a}^{t})),$ (8)
$f_{2}(s^{t},\mathbf{a}^{t})=\sigma(\eta(s^{t},\mathbf{a}^{t})).$ (9)
Here, $f_{1}(\cdot)$ and $f_{2}(\cdot)$ are the same functions in (2), $\eta^{*}$ denotes the optimal predictor neural network weights, while $\eta$ represents the actual predictor weights during training, and $\sigma(\cdot)$ represents the activation function which means the outputs of $f_{1}(\cdot)$ and $f_{2}(\cdot)$ are bounded.
@@
The training samples and the neural network’s weights $\eta$ are bounded, and there exists a set of optimal weights $\eta^{*}$.
@@
If the conditions of Assumption 1 hold, we have:
$\|\eta^{*}-\eta\|<\varepsilon_{w},$ (10)
$\|f_{1}(s^{t},a^{t})-f_{2}(s^{t},a^{t})\|=\|e(t)\|\leq\varepsilon_{e},$ (11)
@@
Given that the weights $\eta$ are bounded, optimal set of weights $\eta^{*}$ exists. This implies that for the bounded sequence of weights in a finite-dimensional vector space, there is a finite maximum distance $\varepsilon_{w}>0$. Due to the boundeness of $\eta$ and the continuity of the neural network function, the neural network outputs for optimal and actual weights does not exceed a bound $\varepsilon_{e}$, as stated in equation (11). ∎
@@
With the deep learning-based predictive model in place, the training error $e(t)$ is uniformly ultimately bounded (UUB).
@@
We construct the Lyapunov function [23] to analyze prediction errors:
@@
$=e^{T}(t)(f_{1}(s^{t},\mathbf{a}^{t})-f_{2}(s^{t},\mathbf{a}^{t})-e(t))$
$=-\|e(t)\|^{2}+\|e(t)\|\|\sigma(\eta^{*}(s^{t},\mathbf{a}^{t}))-\sigma(\eta(s^{t},\mathbf{a}^{t}))\|$
$\leq-\|e(t)\|^{2}+\|e(t)\|\|\eta^{*}(s^{t},\mathbf{a}^{t})-\eta(s^{t},\mathbf{a}^{t})\|$
$=-\|e(t)\|^{2}+\|e(t)\|\varepsilon_{w}$
$=-\|e(t)\|(\|e(t)\|-\varepsilon_{w}).$
@@
When the error norm $\|e(t)\|$ falls short of $\varepsilon_{w}$, the derivative of the Lyapunov function becomes positive, indicating an escalation in the training error up to the point where $\|e(t)\|$ reaches the threshold $\varepsilon_{w}$. If the error norm equals or exceeds $\varepsilon_{w}$, then $\dot{V}(e(t))$ will become non-positive, leading to an error norm that is ultimately bounded by $\varepsilon_{e}$.
In light of these findings, it can be concluded that the training error will asymptotically be less than or equal to $\varepsilon_{e}$, which proves the robustness of the predictor model. ∎
@@ IV-C Integration with MPC
By leveraging the deep learning-based predictor, we integrate it into MPC to cope with the implicit dynamics. The goal of MPC in this work is to find a suitable control signal $\mathbf{a}^{t}$ through online optimization. In this work, given prediction and control horizons $T$, the optimization problem of MPC can be formulated as follows:
@@
$+\sum_{\tau=t+1}^{t+T}\lambda_{\tau}\left(f\left(s^{\tau},\mathbf{a}^{\tau}\right)-\hat{s}^{\tau}\right)$
@@
In SQP, each iteration involves solving a al Quadratic Programming (QP) subproblem to obtain a search direction for the optimization variables. The QP subproblem is formulated based on a quadratic approximation of the objective function and a linear approximation of the constraints around the current iterate $k$. And the QP subproblem for a particular iteration $(s^{k},\mathbf{a}^{k})$ can be stated as follows: